# Improving Relational Database Interactions with Large Language Models: Column Descriptions and Their Impact on Text-to-SQL Performance

### v1

@@ Abstract
Relational databases often suffer from uninformative descriptors of table contents, such as ambiguous columns and hard-to-interpret values, impacting both human users and Text-to-SQL models. This paper explores the use of large language models (LLMs) to generate informative column descriptions as a semantic layer for relational databases. Using the BIRD-Bench development set, we created ColSQL, a dataset with gold-standard column descriptions generated and refined by LLMs and human annotators. We evaluated several instruction-tuned models, finding that GPT-4o and Command R+ excelled in generating high-quality descriptions. Additionally, we applied an LLM-as-a-judge to evaluate model performance. Although this method does not align well with human evaluations, we included it to explore its potential and to identify areas for improvement. More work is needed to improve the reliability of automatic evaluations for this task. We also find that detailed column descriptions significantly improve Text-to-SQL execution accuracy, especially when columns are uninformative. This study establishes LLMs as effective tools for generating detailed metadata, enhancing the usability of relational databases.
@@ 1 Introduction
Figure 1: Example of a question and golden SQL query pair from the BIRD-Bench text-to-SQL benchmark illustrating the need for external knowledge. For a text-to-SQL system to produce a correct SQL query, it needs to know that column A2 and A3 correspond to the district name and region name respectively, which is not intuitively apparent directly from the original column names.
Structured data in relational databases is crucial for modern information systems. It not only stores essential business data but also allows organizations to analyze trends, optimize operations, and make informed decisions, maintaining a competitive edge in today’s fast-paced economy. In this setting, text-to-SQL, which converts natural language queries into SQL commands, enhances access to structured data by reducing the need for expert knowledge.
However, working with these databases often presents challenges due to complex table schemas and vague descriptors used for table contents, such as ambiguous column names and hard-to-interpret values Talaei et al. (2024). These ambiguities can significantly hinder the effectiveness of both human users and generative models employed in text-to-SQL tasks. A good illustration of such an ambiguity can be found in the district table in the financial database of the BIRD-Bench text-to-SQL benchmark Li et al. (2023), as illustrated in Figure 1. The meaning of columns labeled A2 and A3, which represent the district name and its respective region, are not immediately clear from the names alone.
In response to this challenge, the introduction of descriptive column descriptions as a semantic layer over the existing database structures can dramatically improve their usability. However, manually crafting such column descriptions is a time-consuming and labor-intensive process. The large amount of relational databases in use across society today, underscores the need for automated systems that can generate natural language descriptions for SQL databases in order to enhance their interpretability for both human users and text-to-SQL models.
@@
The recent advancements in LLMs have shown that they can produce coherent and contextually relevant text across various domains, making them potential tools for automating such metadata creation [ref]. By leveraging LLMs, the hypothesis is that the labor-intensive process of generating metadata for relational databases can be significantly simplified. Previous research has used generative LLMs to expand abbreviated column names in a SQL setting Zhang et al. (2023), but to the best of our knowledge, no works exist for generating column descriptions in the general case across a wider array of column types and formats.
In this paper we will for this reason explore how well LLMs can generate column descriptions that contain detailed and relevant information. To enable this study, we first create a dataset for column description generation. We use the development set of BIRD-Bench Li et al. (2023) and its databases as the starting point for our dataset. While BIRD-Bench contains some column descriptions, it does not have complete coverage over all columns, and the existing descriptions are often uninformative.
We first improve the existing column descriptions in BIRD-Bench using an LLM to generate more informative descriptions, demonstrating the usefulness of an LLM in a scenario where some initial metadata exists. Secondly, expert human annotators manually correct the generated descriptions to create a refined dataset of gold-standard references as well as annotating the difficulty of generating a column description for each specific column. This corrected dataset serves as a benchmark for evaluating models on column description generation.
We then apply several instruction-tuned models of different sizes and types to our dataset and perform an extensive human evaluation for all models. We find that both GPT-4o and Command R+ frequently generate column descriptions that have perfect overlap with the gold standard.
In addition to human evaluations, we use an LLM-as-a-judge to evaluate model performance, studying how this automatic method correlates with human annotations and its viability for ranking models on this dataset. We find that the LLM judge does not align well with human evaluations and more work is needed to refine the LLM judge for this task.
To assess the general utility of column descriptions, we apply a model to a text-to-SQL task on the BIRD-Bench development set, both with and without column descriptions. Our results demonstrate that column descriptions aid in more accurately generating SQL queries, underscoring their importance in such tasks, especially for columns identified as having limited information. To further verify the importance of column descriptions in scenarios with limited information, we create an alternative version of the BIRD-Bench development set featuring completely uninformative column names. In this extreme scenario, we find that the inclusion of column descriptions is crucial for text-to-SQL generation.
Main Takeaway: Our study demonstrates that LLMs can be a useful tool for generating column descriptions, and including detailed column descriptions leads to more accurate text-to-SQL query generation, particularly in scenarios where column information is limited or uninformative. Future work will focus on applying this methodology to generate other types of metadata, such as value descriptions, to improve the usability of relational databases for humans and generative LLMs.
Database No. of Tables No. of Columns Easy Medium Hard Very Hard
@@
Table 1: Number of tables and columns, and the distribution of column difficulty across the databases in the ColSQL dataset.
2 Related Work
2.1 Table Understanding and Metadata Augmentation Tasks
In recent years, significant effort has been devoted to simplifying the understanding of tabular data. Table-to-text Zhao et al. (2023b, a); Kasner et al. (2023); Yang et al. (2022); Gong et al. (2019) and table question answering Pal et al. (2023); Xie et al. (2022); Herzig et al. (2020) aim at developing models able to understand structured tabular data and natural language questions to perform reasoning and tasks across tables.
More relevant to our work are previous studies focusing on column type detection and column relationship identification for improving the understanding of tables. Column type detection involves classifying columns into predefined semantic categories, such as dates or currencies, to enhance data interpretation Suhara et al. (2022); Zhang et al. (2020). Column relationship identification focuses on determining the structural relationships between columns, such as primary and foreign keys, to understand structured data in tables Wang et al. (2021); Deng et al. (2022). Unlike these approaches, our method of generating column descriptions using large language models aims to produce natural language explanations of column contents, offering a more comprehensive and contextually rich understanding of the data.
While the work of Zhang et al. (2023) focuses on expanding abbreviated column names to improve database schema readability and table understanding tasks using LLMs, our work goes further by generating full descriptions for column contents and considering a much more diverse range of column names.
2.2 Metadata Usage in Text-to-SQL
The shift from traditional query languages to natural language interfaces through text-to-SQL has been significantly advanced by current LLMs. However, the inherent ambiguities and missing information in database schemas necessitate the incorporation of external knowledge and metadata to improve conversion accuracy Zhang et al. (2024). Previous studies have explored various approaches to address this challenge. Talaei et al. (2024) propose a pipeline that retrieves relevant data and context from a knowledge store to generate accurate SQL queries, leveraging a data store based on the original BIRD-Bench metadata files. Similarly, Dou et al. (2022) enhance the text-to-SQL process by using external domain-specific knowledge, such as mathematical formulas and publicly available resources.
Sun et al. (2024) introduce question-specific database content to improve tuning performance, while Hong et al. (2024) utilize a tailored data expert LLM to provide useful knowledge to the SQL-generating LLM. Petrovski et al. Petrovski et al. (2018) take a different approach by vectorizing entire data columns into embeddings, thus eliminating the need for traditional descriptions and metadata in the text-to-SQL process. These studies collectively validate that incorporating external knowledge and metadata can significantly enhance text-to-SQL systems.
3 Column Description Dataset Creation
To enable our study on generating informative column descriptions, we create ColSQL, a new dataset based on the development set of the widely-used text-to-SQL benchmark, BIRD-Bench Li et al. (2023) [^1] BIRD-Bench is licensed under $C$ BY-SA 4.0, permitting transformation and reuse of the dataset.. This dataset consists of 799 columns across 11 different databases and domains, which are presented in Table 1. We selected BIRD-Bench as the basis for our dataset because it allows us to test column description generation across a broad set of domains. Additionally, the natural language queries in BIRD-Bench were designed so that external knowledge (e.g., metadata) is necessary to perform well on the task. In contrast, the WikiSQL Zhong et al. (2017) and SPIDER Yu et al. (2018) text-to-SQL benchmarks primarily uses exact column names in its natural language queries, making them unsuitable for our use case.
While BIRD-Bench already includes column descriptions, we discovered that many columns either lacked descriptions entirely or contained insufficient or erroneous information. This necessitated improving the existing column descriptions to create an adequate gold standard for evaluating models.
In the following sections, we first define what constitutes a good column description. We then describe how we improved BIRD-Bench’s descriptions using generative LLMs and human annotators. Additionally, we explain how we annotated the difficulty level for generating a description from each column.
3.1 What is a good column description?
Natural language descriptions are crucial for addressing the often subtle ambiguities and complexities found in column names. Previous works that relied on a single word or type to describe columns are insufficient for capturing the full context and meaning. A more detailed description ensures that the interpretation of the column is free of ambiguity, facilitating better understanding and more accurate querying.
Defining exactly what constitutes a good description in a set of rules is a complex task, because of the complex nature of natural language and the complexities in SQL databases. However, we used the following criteria to define a good description:
  • It should eliminate ambiguity and enhance the interpretability of a column.
  • It should provide clear and precise information about the column, specifying the nature of the data it holds and its context within a table.
  • It should include information about the subject of the reference, e.g., instead of "The name," use "The name of the client that made the transaction" to ensure clarity.
  • It must be a complete and grammatically correct sentence that explicitly describes what the column represents.
We do not require the inclusion of interpretations of specific values inside the database, leaving the interpretation and generation of metadata for database values for future work.
Given the numerous ambiguities and complexities in SQL databases due to omitted information, we invite the community to engage in discussions about what constitutes a good column description and the necessary metadata to include.
3.2 Improving Existing Column Descriptions with LLMs
Label Original LLM Updated
Perfect 40.24% 70.80%
Somewhat correct 38.80% 18.80%
Incorrect 1.88% 0.06%
I can’t tell 0.81% 7.27%
No description 18.27% 3.07%
Table 2: Quality of the original BIRD-Bench column descriptions and the LLM improved descriptions.
We aimed to improve existing column descriptions using LLMs for two reasons: as a preliminary step before human annotators create the gold standard dataset, and to evaluate the effectiveness of LLMs in improving existing column descriptions.
To improve column descriptions, we used GPT-4o, providing it with the BIRD-Bench column descriptions, corresponding schema in CREATE_TABLE statements, example data rows from the corresponding table, and instructions to generate descriptions for given columns. The model was given clear guidelines and in-context examples of the desired description format. The full prompt template is in Appendix B.4.
Two human annotators who are authors of this paper, fluent in English, and experts in SQL, assessed the original BIRD-Bench descriptions and the LLM-generated descriptions, independently assigning each description one of the following labels:
  • Perfect: Contains enough information so that the interpretation of the column is completely free of ambiguity.
  • Somewhat Correct: Matching the perfect description but verbose with redundant information (although without any incorrect or misleading information).
  • Incorrect: Contains incorrect information.
  • No Description: Description is missing.
  • I can’t tell: Annotator cannot determine the description’s accuracy based on available data.
The full annotator guidelines and more detailed descriptions of the labels are presented in Apendix B.1.2.
Annotators’ assessments of the two datasets are shown in Table 2. There was a substantial level of agreement between annotators, with a Cohen’s kappa of $0.680.68$ on the original BIRD-Bench descriptions and $0.610.61$ on the improved generated descriptions. The GPT-4o descriptions showed a significant improvement over the original BIRD-Bench descriptions. While not perfect, LLM-generated descriptions can be a useful tool for improving existing column metadata when there are already quality issues in the existing column metadata.
Figure 2: Mean quality scores for the generated column descriptions from each model on the ColSQL dataset, as rated by the human evaluators.

### v3

# Synthetic SQL Column Descriptions and Their Impact on Text-to-SQL Performance

@@ Abstract
Relational databases often suffer from uninformative descriptors of table contents, such as ambiguous columns and hard-to-interpret values, impacting both human users and text-to-SQL models. In this paper, we explore the use of large language models (LLMs) to automatically generate detailed natural language descriptions for SQL database columns, aiming to improve text-to-SQL performance and automate metadata creation. We create a dataset of gold column descriptions based on the BIRD-Bench benchmark, manually refining its column descriptions and creating a taxonomy for categorizing column difficulty. Through evaluating several LLMs, we find that incorporating these column descriptions consistently enhances text-to-SQL model performance, particularly for larger models like GPT-4o, Qwen2 72B and Mixtral 22Bx8. However, models struggle with columns that exhibit inherent ambiguity, highlighting the need for manual expert input. Notably, Qwen2-generated descriptions, containing by annotators deemed superfluous information, outperform manually curated gold descriptions, suggesting that models benefit from more detailed metadata than humans expect. Future work will investigate the specific features of these high-performing descriptions and explore other types of metadata, such as numerical reasoning and synonyms, to further improve text-to-SQL systems. The dataset, annotations and code will all be made available.
@@ 1 Introduction
Text-to-SQL, which involves converting natural language queries into Structured Query Language (SQL), holds promise as it facilitates broader access to structured data in SQL databases by reducing the need for expert knowledge. However, working with SQL databases often presents challenges due to ambiguous or uninformative column names, hindering the effectiveness of both human users and generative models employed in text-to-SQL tasks [1, 2]. An illustration of such uninformative column names can be found in the district table in the financial database of the BIRD-Bench text-to-SQL benchmark [3], which is shown in Figure 1. The columns labeled A2-A16 have completely uninformative names and are therefore hard to interpret without either amassed domain knowledge or access to external database documentation.
To address this challenge, recent works have explored using natural language descriptions or metadata as a semantic layer atop existing database structures, significantly enhancing their usability [2, 4, 3]. The critical role of such metadata is further underscored by the fact that all of the top 20 models on the BIRD-Bench benchmark incorporate by the benchmark-provided metadata to attain their high scores, as evidenced by the Oracle Knowledge checkbox on the benchmark’s website [^1] https://bird-bench.github.io/. However, this metadata was manually crafted, making it both time-consuming and labor-intensive to produce. Given the widespread use of relational databases across society, each varying in structure and size, there is a pressing need for automated systems capable of generating natural language descriptions for SQL databases to enhance their interpretability for both human users and text-to-SQL models.
A related area to generating column descriptions is semantic column type detection, where models infer a column’s data type based on its content or schema [5, 6]. However, these methods typically provide only basic type annotations, lacking the context needed to fully understand the column’s purpose. For example, identifying a column as "dates" doesn’t clarify whether they are transaction or birth dates. This limited detail leaves users and text-to-sql models to interpret the column’s exact meaning. What’s needed are systems that generate detailed, human-readable descriptions of the column’s meaning.
@@
The recent advancements in LLMs have shown that they can produce coherent and contextually relevant text across various domains [7], making them potential tools for automating such metadata creation. In the context of SQL databases, the hypothesis is that LLMs could leverage multiple signals from a database to generate meaningful column descriptions, including the schema structure, relationships between tables, and the values present within the database itself. By analyzing these signals, LLMs can go beyond simple type detection, attempting to infer a column’s broader purpose and a more detailed explanation. Previous research has used generative LLMs to expand abbreviated column names in a SQL setting [8], but to the best of our knowledge, no works exist for generating more detailed descriptions across a wider array of column types and formats.
Motivated by these observations, we therefore investigate the capability of LLMs to generate column descriptions that are both detailed and contextually relevant, and how such metadata can enhance the effectiveness and robustness of text-to-SQL models. To enable this study, we first construct a dataset specifically for database description generation using a mix of manual and automated techniques to improve the existing column descriptions provided by the BIRD-Bench benchmark. Second, expert human annotators annotate the difficulty of generating a column description for each specific column to provide a deeper understanding of what table contexts are possible for an LLM to understand.
We then evaluate several instruction-tuned models for column description generation and assess the impact of different types of column descriptions on text-to-SQL model performance. Our findings reveal that models such as Mixtral 22Bx8 and Command R+ excel in generating high-quality descriptions. However, generating accurate descriptions for columns with inherent ambiguity remains a challenge. Despite this, our experiments show that including detailed column descriptions, whether manually refined or synthetically generated, consistently enhances text-to-SQL accuracy across a range of LLMs.
Figure 1: Example rows from the district table in the BIRD-Bench benchmark dataset. This table illustrates a typical schema with uninformative column names such as A2-A16, making interpretation difficult without external documentation or domain knowledge. The column names provide little semantic meaning, and the accompanying data varying amount of information, complicating their use in text-to-SQL query generation and requiring additional metadata for effective database interactions.
In this paper, we make the following contributions:
  • We create a new dataset based on the BIRD-Bench benchmark, improving the quality of column descriptions through a combination of LLM-generated outputs and manual refinements. This dataset provides a strong foundation for evaluating column description generation.
  • We introduce a difficulty annotation scheme to classify columns based on the ease with which meaningful descriptions can be generated, offering insights into the inherent challenges faced by models.
  • We evaluate several state-of-the-art LLMs, including GPT-4o, Command R+, Qwen2 72B, and others, in their ability to generate detailed and contextually accurate column descriptions, assessing their performance against the gold standard.
  • We conduct a series of text-to-SQL evaluations to analyze the effect of different types of column descriptions on SQL query generation, revealing that more detailed metadata, even if perceived as superfluous by humans, can significantly improve model performance.
  • Finally, we propose a two-step process for handling inherently ambiguous columns, combining automated generation with human expert intervention. We also highlight key areas for future research, including exploring additional metadata types to enhance text-to-SQL systems further.
2 Related Work
2.1 Table Understanding and Metadata Augmentation Tasks
In recent years, significant effort has been devoted to simplifying the understanding of tabular data. Table-to-text [9, 10, 11, 12, 13] and table question answering [14, 15, 16] aim at developing models able to understand structured tabular data and natural language questions to perform reasoning and tasks across tables.
Previous studies on column type detection and relationship identification have aimed to enhance table understanding by classifying columns into semantic categories (e.g., dates, currencies) [5, 6] and identifying structural relationships (e.g., primary and foreign keys) [17, 18]. In contrast, our approach uses LLMs to generate natural language descriptions, providing a more comprehensive and contextually rich understanding of column contents.
While the work of Zhang et al. [8] focuses on expanding abbreviated column names to improve database schema readability and table understanding tasks using LLMs, our work goes further by generating full descriptions for column contents and considering a much more diverse range of column names.
2.2 Metadata Usage in Text-to-SQL
The shift from traditional query languages to natural language interfaces through text-to-SQL has been significantly advanced by current LLMs. However, the inherent ambiguities and missing information in database schemas necessitate the incorporation of external knowledge and metadata to improve conversion accuracy [19]. Previous studies have explored various approaches to address this challenge. Li et al. [2] propose a pipeline that retrieves relevant data and context from a knowledge store to generate accurate SQL queries, leveraging a data store based on the original BIRD-Bench metadata files. Similarly, Dou et al. [20] enhance the text-to-SQL process by using external domain-specific knowledge, such as mathematical formulas and publicly available resources.
Sun et al. [21] improve tuning performance using question-specific database content, while Hong et al. [22] employ a data expert LLM to aid SQL generation. Petrovski et al. [23] take a different approach by embedding entire data columns, removing the need for traditional metadata. These studies demonstrate the value of external knowledge and metadata in enhancing text-to-SQL systems.
3 Column Description Dataset Creation
To enable our study on generating informative column descriptions, we create a new dataset based on the development set of the widely-used text-to-SQL benchmark, BIRD-Bench [3] [^2] BIRD-Bench is licensed under CC BY-SA 4.0, permitting transformation and reuse of the dataset.. This dataset consists of 799 columns across 11 different databases and domains, which are presented in Table 1. We selected BIRD-Bench as the basis for our dataset because it allows us to test column description generation across a broad set of domains. Additionally, the natural language queries in BIRD-Bench were designed so that external knowledge (e.g., metadata) is necessary to perform well on the task. In contrast, the WikiSQL [24] and SPIDER [25] text-to-SQL benchmarks primarily uses exact column names in its natural language queries, making them unsuitable for our use case.
While BIRD-Bench already includes column descriptions, many lacked sufficient detail or contained inaccuracies. To address this, we improved the existing descriptions using a combination of automated LLM and manual refinements, which allowed us to create a gold standard for evaluating model performance. Note that we chose to focus on generating column descriptions for this work, and leave the generation of descriptions of values inside the database for future work.
3.1 Improving Existing Column Descriptions with LLMs
We aimed to improve existing column descriptions using LLMs both as a preliminary step before human annotators created the gold standard dataset, and to evaluate the effectiveness of LLMs in improving descriptions.
To improve existing column descriptions, we provided GPT-4o the BIRD-Bench column descriptions, corresponding schema, example data and an instruction to generate new descriptions. The full prompt template can be found in Appendix B.9. The original BIRD-Bench descriptions and the generated ones were then independently assessed by two human annotators experts in SQL. Descriptions were categorized as either "Perfect," "Somewhat Correct," "Incorrect," "No Description," or "I can’t tell." The full annotator guidelines and more detailed descriptions of the labels are presented in Apendix B.7.2.
The results show that the original BIRD-Bench descriptions were "Perfect" for 40.24% of columns, while LLM-generated descriptions improved this to 70.80%. The percentage of "Somewhat Correct" descriptions decreased from 38.80% to 18.80% after LLM improvement, and errors were nearly eliminated (dropping from 1.88% to 0.06%). A table showcasing the full results of the annotations can be found in Appendix B.6.
Cohen’s kappa indicated substantial agreement between annotators, with scores of 0.68 for the original descriptions and 0.61 for the LLM-generated ones. While not perfect, LLM-generated descriptions significantly improved over the original dataset and proved useful for improving the quality of column descriptions with existing issues. Full details of the results are presented in Appendix B.6.
3.2 Creating the Gold Standard Dataset
Since not all LLM-generated improvements were perfect, manual corrections of the generations were necessary to achieve the high standard required for a gold standard dataset. First, disagreements in scores were addressed by reviewing each other’s annotations and then through discussion to reach a consensus. Descriptions that did not achieve a “Perfect” quality score from both annotators after this phase were manually corrected until both annotators agreed on a “Perfect” rating. The final set of descriptions was then reviewed to ensure they provided clear, comprehensive, and contextually rich information for each column. Through this process, we believe we have created a high quality dataset serving as a reliable reference for evaluating the performance of models in generating column descriptions. The final composition of the dataset can be found in Table 1. Examples of gold descriptions can be found in Appendix B.8.
Database # tables # columns Self-Evident Context-Aided Ambiguity-Prone Domain-Dependent
@@ 3.2 Creating the Gold Standard Dataset
Table 1: Number of tables and columns, and the distribution of column difficulty across the databases in our column description dataset.
3.3 Annotating Column Difficulty
After finalizing the gold standard descriptions, columns were categorized into four difficulty levels: "Self-Evident," "Context-Aided," "Ambiguity-Prone," and "Domain-Dependent," based on the amount of information available in the database to generate accurate descriptions.
For "Self-Evident" columns, like birth_date in the Client table, the schema alone was sufficient to generate an accurate description ie. the column’s meaning was self-evident. "Context-Aided" columns, such as short_state in the zip_code table, required signals from both the schema and the data for accurate interpretation. This suggests that while some additional context is required, the column’s role can still be deduced by looking at the data inside the database. "Ambiguity-Prone" columns posed challenges even with the available schema and data due to uncertainties; for example, in the formula_1 database’s results table, the rank column has data indicating rankings, but without context, it is hard to determine if it refers to race position, qualifying rank, or another metric. The description, "The overall rank of the driver’s fastest lap time in the race," clarifies its purpose. "Domain-Dependent" columns, like A11 in the districts table, were impossible to describe without access to additional documentation or domain knowledge.
Two independent annotators, separate from those who created the gold descriptions, assessed the difficulty, achieving substaintial agreement (Cohen’s kappa of 0.61). Disagreements were resolved through discussion, resulting in a final difficulty score for each column. Most columns fell into the "Self-Evident" and "Context-Aided" categories, while "Domain-Dependent" columns were primarily found in the Financial domain.
4 Experiments
4.1 Generating Column Descriptions
To evaluate the effectiveness of LLMs in generating column descriptions, we used our dataset presented in Section 3.2. We evaluated the following models: GPT-4o, Mixtral-8x22B-Instruct-v0.1 [26], Command R+, Qwen2-72B-Instruct [27], Codestral-22B-v0.1, and Mistral-7B-Instruct-v0.3 [28]. These models were selected to provide a diverse range of perspectives, including a state-of-the-art closed-source model (GPT-4o), several newly released high-performing open-source models (Qwen2-72B, Mixtral-8x22B, Codestral-22B, Command R+), and a smaller model (Mistral-7B).
Inference for the open-source models was performed on several A100 GPUs, using BitsandBytes [29] to load the models in 8-bit format. We set the model temperature to 0.7.
4.2 Identifying Ambiguous Columns
For SQL databases, some columns might be inherently ambiguous as described in Section 3.3, making it difficult to generate accurate descriptions without domain knowledge. To address this, we conducted a small ablation study to evaluate whether LLMs could autonomously recognize the "Ambiguity-Prone" or "Domain-Dependent" columns—specifically those that exhibit inherent ambiguity. The goal was to assess the model’s ability to detect situations where the schema and data alone are insufficient for making unambiguous predictions about the columns’ contexts.
In our experiment, we modified the prompt provided to the LLMs, including the instruction: "If the details in the schema do not suffice to ascertain what the columns meaning is without ambiguity, return: ’Not enough information to make a valid prediction.’" This addition was designed to prompt the LLM to identify and flag ambiguous columns rather than generating potentially misleading descriptions.
By leveraging this capability, we propose a two-step approach: first, automatically generating descriptions for columns the LLMs identify as unambiguous, and second, involving human experts to annotate the remaining columns where inherent ambiguity is detected.
4.3 Text-to-SQL Evaluations
To study the usefulness of column descriptions for text-to-SQL systems, we first evaluated the same LLMs as in the description generation step in a zero-shot setting on the development set of BIRD-Bench, which contains 1534 natural language questions and SQL query pairs. Our goal was to isolate and understand the effects of column descriptions on model performance.
The experimental setup involved five scenarios: (1) Without any column descriptions, where the model operates with only the database schema; (2) With BIRD-Bench original column descriptions, utilizing the original descriptions provided by BIRD-Bench in addition to the schema; (3) With gold standard descriptions, where the model uses our manually refined gold standard descriptions in addition to the schema; (4) With GPT-4o generated column descriptions, where the model uses descriptions generated by the best model from Section 4.1 in addition to the schema; and (5) With Qwen2 generated descriptions, where the model uses descriptions rated with a high amount of Superflous information. We included both the GPT-4o and the Qwen2-generated descriptions to test whether our humanly defined categories of "Perfect" and "Superfluous information" are aligned with what is practically useful for LLMs in the text-to-SQL task.
In all scenarios, we included only column descriptions as the available metadata alongside the database schema. We set the model temperature to 0. The full prompt for the text-to-SQL generation is presented in Appendix B.9.
5 Results
Figure 2: Visual representation of LLM performance in generating SQL column descriptions.